OpenStudy (anonymous):

-7x + 3 = -15/2 Solving Equations Containing Fractions

OpenStudy (anonymous):

subtract 3, divide by \(-7\)

OpenStudy (anonymous):

\[-7x+3=-\frac{15}{2}\] \[-7x=-\frac{15}{2}-3=-\frac{21}{2}\] \[x=-\frac{21}{2}\times \frac{1}{-7}=\frac{3}{2}\]

OpenStudy (anonymous):

why did you subtract 3

OpenStudy (anonymous):

because you had \(-7x+3\) on one side of the equal sign, and you wanted to get \(x\) by itself

OpenStudy (anonymous):

in order to get rid of the \(3\) you have to subtract it from both sides like for example if you had \[x+3=8\] then you know \[x=5\] becaues \(5+3=8\) or because \(8-3=5\)